OpenStudy (klepez):

how do you solve (x/x-4 + 6/x+4)/ 3/2x^2-32

OpenStudy (reemii):

it's \((x/x-4 + 6/x+4)/ (3/2x^2-32)\) right?

OpenStudy (klepez):

yes

OpenStudy (reemii):

if it's "solve", then there should be an equality sign, something like \((x/x−4+6/x+4)/(3/2x^2−32) = 0\) (or =1). but even without that, you can do something aout this fraction. 1) work on the numerator: you must make one fraction (not a sum of two) 2) factorize all the expressions and make simplifications.

OpenStudy (reemii):

objective: to write your fraction as \(\frac{\frac{a}{b}}{\frac{c}{d}}\). maybe \(a\) and \(c\) have common factors, maybe \(b\) and \(d\) have common factors.

OpenStudy (reemii):

It starts like: \((\frac{x}{x-4} + \frac{6}{x+4}) / (3/(2x^2-32)) = \frac{x(x+4)+6(x-4)}{(x-4)(x+4)} / (3/(2x^2-32)) = \frac{x(x+4)+6(x-4)}{(x-4)(x+4)} \frac{2x^2-32}{3}\). when you factorize \(2x^2-32\) you'll see "something".

OpenStudy (reemii):

oops. i wrote \[ \frac{x(x+4)+6(x-4)}{(x-4)(x+4)} \frac{2x^2-32}{3} \]on the right.

OpenStudy (reemii):

do you get something interesting?

OpenStudy (klepez):

no not really..

OpenStudy (reemii):

\(2x^2-32 = 2(x^2 - 16) = 2(x+4)(x-4)\) (the last step used the formula for a difference of two squared numbers: \(a^2-b^2=(a-b)(a+b)\).) here you got some factors that will cancel out other factors already present in the fraction. see that? :D

OpenStudy (klepez):

ya

OpenStudy (reemii):

write your final answer here, and can you tell why the question is "solve" ? im puzzled

OpenStudy (klepez):

its not it just needs to be awnsered

OpenStudy (reemii):

\(\frac{x^2+10x-24}{\text{cancels out}}\frac{2(\text{cancels out})}{3} = \frac23(x^2+10x-24)\). did you find that..?

OpenStudy (klepez):

ya

OpenStudy (reemii):

well done

OpenStudy (klepez):

is that the final awnser?

OpenStudy (reemii):

There's nothing more to do (except doing \(\frac23x^2 + \frac2310x - \frac23 ....\) but that's not interesting). Yes it's the final answer.

OpenStudy (klepez):

ok thank you for the help!